If x – (7/8) = 7/8, then x =                             
 If x – (7/8) = 7/8, then x =                             

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider the equation, x–(7/8) = 7/8
Transpose, -7/8 to RHS then it becomes 7/8.
Then, x = (7/8) + (7/8)
x = (14/8) [divide by 2]
x = 7/4

Thanks